### v1

# No Images, No Problem: Retaining Knowledge in Continual VQA with Questions-Only Memory

Abstract
Continual Learning in Visual Question Answering (VQACL) requires models to learn new visual-linguistic tasks (plasticity) while retaining knowledge from previous tasks (stability). The multimodal nature of VQACL presents unique challenges, requiring models to balance stability across visual and textual domains while maintaining plasticity to adapt to novel objects and reasoning tasks. Existing methods, predominantly designed for unimodal tasks, often struggle to balance these demands effectively. In this work, we introduce QUestion-only replay with Attention Distillation (QUAD), a novel approach for VQACL that leverages only past task questions for regularization, eliminating the need to store visual data and addressing both memory and privacy concerns. QUAD achieves stability by introducing a Question-only Replay mechanism that selectively uses questions from previous tasks to prevent overfitting to the current task’s answer space, thereby mitigating the out-of-answer-set problem. Complementing this, we propose Attention Consistency Distillation, which uniquely enforces both intra-modal and inter-modal attention consistency across tasks, preserving essential visual-linguistic associations. Extensive experiments on VQAv2 and NExT-QA demonstrate that QUAD significantly outperforms state-of-the-art methods, achieving robust performance in continual VQA [^1] Work done during an internship at Computer Vision Center (CVC), Universitat Autonoma de Barcelona.. Code is available at: https://github.com/IemProg/QUAD
1 Introduction
Continual learning (CL) seeks to enable models to incorporate new information while retaining previously learnt knowledge, thereby addressing the problem of catastrophic forgetting (CF) [48, 49]. This capability is essential in dynamic real-world settings, where models must continuously adapt to evolving data while preserving previously acquired knowledge [85]. While significant advances have been made in CL, research has predominantly focused on unimodal tasks, particularly image classification [70, 69, 33, 22, 47, 87]. However, real-world applications often demand more sophisticated multimodal learning approaches, especially for tasks that integrate visual and textual information. Visual Question Answering (VQA) stands as a representative multimodal task, requiring models to simultaneously process and reason over both images and natural language. For instance, answering questions like “What color is the car on the left?” or “How many trucks are visible?” demands both object recognition from the image and an understanding of the linguistic query structure [85, 60, 78, 8, 72].
Figure 1: Comparison of continual learning methods for Visual Question Answering (VQA) in terms of storage, forgetting, and privacy. Memory-free methods ensure privacy but suffer from high forgetting. Memory-rehearsal methods reduce forgetting but raise privacy concerns by storing sensitive images (e.g. people’s identities, car plates). Our approach QUAD, only stores questions and avoids image storage, preserving privacy while achieving low forgetting by leveraging question-based regularisation to effectively solve the out-of-answer set problem.
The emerging field of Visual Question Answering Continual Learning (VQACL) [85] presents unique challenges at the intersection of continual learning and multimodal reasoning. VQACL requires maintaining a balance between stability—the preservation of past knowledge—and plasticity—the ability to learn new information—across both visual and linguistic modalities [85, 3]. This dual-modality requirement, combined with the need for generalisation ability, introduces additional complexity in learning dynamics. Models must not only retain their understanding of previously encountered visual elements and language patterns but also generalise to unseen objects and novel question formats. For example, a model that masters counting vehicles should be able to transfer this counting capability to novel object categories, such as bicycles [23].
Memory-based methods in CL [7, 6, 67, 85] effectively mitigate catastrophic forgetting through storing and replaying exemplar data. In the VQACL setting, these approaches store complete image-question pairs, necessitating significant memory resources—often thousands per task (e.g., 5000 samples in VQACL [85]). This extensive storage of visual data raises two critical challenges. First, it imposes substantial computational and storage burdens, which can be infeasible in resource-constrained real-world applications. Second, and more importantly, it raises profound privacy concerns. Visual data often contains sensitive and personally identifiable information, especially in domains like healthcare, finance, and surveillance [64, 42], where stringent regulations such as GDPR [19, 81] govern data usage and storage. In contrast, textual data (the questions) is typically non-identifiable and generic, inherently posing less privacy risk. Memory-free methods [35, 1, 83, 33], though eliminating storage and privacy concerns by avoiding data retention entirely, exhibit suboptimal performance in multimodal settings like VQACL. This dichotomy between memory-rehearsal and memory-free approaches raises an intriguing questions: Is it truly necessary to store visual data? Could storing only past questions suffice to mitigate forgetting? To explore these questions, we propose a novel intermediate setting: VQACL with Question-only Rehearsal (VQACL-QR) (Fig. 1). In this framework, only questions from past tasks are stored, eliminating the need for visual data storage. It provides a practical, privacy-preserving solution for continual VQA setting.
To address VQACL-QR , we propose QUestion-only replay with Attention Distillation (QUAD), a novel distillation method to balance stability and plasticity using only past questions. Our approach introduces two key contributions, each targeting specific challenges in continual VQA. First, we introduce a Question-only Replay mechanism, which utilises only questions from past tasks to regularise the current model with a strategic question selection process that directly tackles the out-of-answer-set problem – model tend to overfit to the current task’s answer space, leading to incorrect responses for previous tasks’ questions (see Fig. 1). Second, we propose attention consistency distillation, a novel attention distillation strategy that preserves the model’s attention patterns across tasks. Our distillation strategy enforces both intra-modal consistency (language-to-language, visual-to-visual) and inter-modal consistency (language-to-visual), ensuring that the model consistently focuses on relevant regions while adapting to new information.
Through extensive experiments on standard VQACL benchmarks (VQAv2, and NExT-QA datasets), we demonstrate that QUAD achieves state-of-the-art performance, surpassing both memory-free approaches and traditional rehearsal methods that store complete image-question pairs. Our findings suggest that storing only questions is sufficient to mitigate forgetting, validating the practicality of the Question-only Rehearsal (VQACL-QR) setting.
2 Related work
Visual Question Answering (VQA) involves answering natural language questions by interpreting visual content [3, 51, 38]. Recent approaches leverage vision-language models (VLMs) built on transformer architectures [16, 61, 76, 37, 8] alongside pre-trained language models [78, 80]. For instance, Cho et al. [9] proposed a generative transformer for VQA that analyses visual elements in conjunction with textual questions. Many methods aim to improve generalisation capabilities by enhancing compositionality, which is essential for cognitive reasoning [30, 36]. For instance, Johnson et al. [28] explored the composition of visual attributes by creating a dataset designed for compositional reasoning, while Whitehead et al. [73] used contrastive learning to enhance compositionality, disentangling reasoning skills from visual concepts. Although these approaches make strides toward compositionality, the implicit decomposition may impact generalisation, and constructing effective contrastive samples remains complex.
Continual Learning (CL) aims to develop frameworks capable of incrementally learning from sequentially arriving datasets. This is a fundamental challenge for many deep learning methods due to catastrophic forgetting (CF) [48]. CL methods can be divided into different categories. While knowledge distillation-based methods constrain mappings between incremental models to prevent forgetting [39, 57, 27, 46, 13, 29], optimisation-based techniques modify the gradient updates during training to minimise interference with previously learnt tasks to mitigate forgetting [44, 5, 59, 79], and representation-based methods focus on learning robust and adaptable feature representations [18, 17, 15, 14]. Recently we have also witnessed the uprisal of prompt-based methods like [71, 70, 68, 62], which utilise visual prompts [43] with pre-trained transformers for CL scenarios. Most of these methods target Class-Incremental Learning (CIL) in unimodal settings; as such, they cannot simultaneously tackle multimodal continuous data and ignore the essential composition generalisation issue for VQA [85, 52]. Our work falls within the distillation-based category, as QUAD enables distillation with only past questions to reduce CF for continual VQA.
Continual VQA. Recent research has increasingly focused on multimodal CL in the context of VQA [10, 24, 52, 63]. While studies like Greco et al. [24] have examined forgetting dynamics in VQA, they often focus on a limited set of tasks and overlook the potential of pre-trained models to mitigate forgetting [50]. The VQACL benchmark [85] evaluated conventional continual learning approaches such as  [7, 6, 67, 85]. Some works have explored specific aspects of continual VQA, such as question types [24], domain adaptation [86], and compositionality [85]. Our work introduces a distillation method for continual VQA while remaining memory-efficient, privacy-conscious, and without any stored prototypes as in the VQACL method [^2] VQACL represents both the setting and the approach for continual learning in Visual Question Answering, as described in  [85]. [85].
3 Question-only replay with Attention Distillation (QUAD)
3.1 Setting Overview
The VQACL setting [85] is designed to assess the capability of a model to adapt to a sequence of tasks, each involving both visual and linguistic inputs, in a continual learning environment. It approaches VQA as a generative task, where the objective is to generate textual answers given an image and a corresponding question [20, 85]. The model encounters a non-stationary stream of data, requiring it to learn and adapt incrementally over time without revisiting prior data. We consider a sequence of $T$ tasks, denoted as $\mathcal{T}^{1},\mathcal{T}^{2},\ldots,\mathcal{T}^{T}$. Each task $\mathcal{T}^{t}$ is characterized by a set of image-question-answer triplets $(x^{t},q^{t},y^{t})$, where $x^{t}\in\mathcal{X}^{t}$ denotes the image, $q^{t}\in\mathcal{Q}^{t}$ represents the question, and $y^{t}\in\mathcal{Y}^{t}$ corresponds to the answer. [^3] The sample index is omitted for clarity. The challenge is to train a model $\phi$ that can effectively learn the current task $\mathcal{T}^{t}$ while retaining the knowledge from all previous tasks $\{\mathcal{T}^{1},\mathcal{T}^{2},\ldots,\mathcal{T}^{t-1}\}$.
In VQACL, the sequence of tasks is organized as a series of $L$ macro-tasks, each comprising $K$ sub-tasks, resulting in a total of $T=L\times K$ tasks. Each macro-task is designed to develop specific reasoning skills such as counting, color identification, or object recognition (i.e linguistic task). For example, in a counting task, the model primarily engages with questions like “How many objects are there?” or “What number is shown?”.
Each linguistic macro-task is further divided into visually-driven sub-tasks. Formally, each macro-task $\mathcal{T}^{t}$ is split into $K$ visually-driven sub-tasks, $\{\mathcal{S}^{t}_{1},\mathcal{S}^{t}_{2},\ldots,\mathcal{S}^{t}_{K}\}$, which are learned sequentially. These sub-tasks $\mathcal{S}^{t}_{k}$ are constructed by grouping the $C$ distinct visual object categories, $\{c_{i}\}_{i=1}^{C}$, into $K$ sets. This hierarchical structure mirrors the continuous nature of the visual and linguistic data streams that the model processes. The VQACL setting introduces two unique challenges for continual learning models. 1) Knowledge Retention: as the model progresses through the sequence of tasks, it must retain knowledge from earlier tasks to perform well on future tasks, where both visual and linguistic modalities must be preserved. 2) Generalization to Novel Compositions: this setting also evaluates the model’s ability to generalize to novel combinations of visual concepts and reasoning skills that it has not encountered during training. This aspect is crucial for real-world applications where new object-skill combinations are frequently encountered. Details of task sequences, object groupings, and novel composition testing is in the Appendix.
Figure 2: Overview of QUAD. Our method consists of three key components designed to maintain stability and plasticity in the VQACL setting. (1) Questions-Only Memory ($\mathcal{M}$) stores questions from past tasks, without visual data. (2) Pseudo-label Distillation ($\mathcal{L}_{\text{PL}}$) leverages answers generated by the previous model $\theta^{t-1}$ for new image-question pairs, guiding the current model $\theta^{t}$ to retain past knowledge, and mitigate the out-of-answer-set problem. (3) Attention Consistency Distillation ($\mathcal{L}_{\text{Att}}$) aligns the self-attention maps between $\theta^{t-1}$ and $\theta^{t}$ to maintain focus on relevant visual-linguistic relationships. The task-specific loss ($\mathcal{L}_{\text{CE}}$) is applied solely to current task samples, promoting adaptation to new data.
3.2 Overview
We propose QUAD, a novel approach for VQACL-QR that eliminates the need to store images from previous tasks by using only past task questions (see Fig. 2). Following several works in continual learning [40, 11], we adopt a regularisation framework where the total learning objective $\mathcal{L}_{\text{VQACL}}$ consists of two main components:
$\mathcal{L}_{\text{VQACL}}=\mathcal{L}_{\text{Plasticity}}+\lambda\mathcal{L}_{\text{Stability}}.$ (1)
The first, the plasticity term $\mathcal{L}_{\text{Plasticity}}$, guides the model’s adaptation to the current task $\mathcal{T}^{t}$, and $\lambda>0$ is a weighting coefficient adjusting the balance between plasticity and stability. Following common practice in VQA [85, 56, 2], we implement this loss using cross-entropy to compare the network’s prediction for an input image-question pair $(x^{t},q^{t})$ with the corresponding annotated answer $y^{t}$:
$\mathcal{L}_{\text{Plasticity}}=\mathbb{E}_{(x^{t},q^{t},y^{t})\sim\mathcal{T}^{t}}\mathcal{L}_{\text{CE}}\left[\phi(x^{t},q^{t}),y^{t}\right].$ (2)
The second loss component, the stability term $\mathcal{L}_{\text{Stability}}$, helps prevent forgetting. In the standard VQACL setting, where images from past tasks can be stored in the memory $\mathcal{M}$, this term can be implemented similarly to the plasticity term, using cross-entropy and averaging the loss over triplets ${}{(x^{m},q^{m},y^{m})\in\mathcal{M}}$. In the VQACL-QR setting, storage of images $x^{m}$ from past tasks in the memory $\mathcal{M}$ is prohibited. To address this limitation, we propose a novel $\mathcal{L}_{\text{Stability}}$ loss tailored for VQACL-QR. This loss combines two complementary strategies—pseudo-labeling $\mathcal{L}_{\text{PL}}$ and attention distillation $\mathcal{L}_{\text{Att}}$— to effectively compensate for the absence of past task images, ensuring robust knowledge retention while adhering to the constraints of the VQACL-QR framework. In our case, the stability term can be written as:
$\mathcal{L}_{\text{Stability}}=\mathcal{L}_{\text{PL}}+\mathcal{L}_{\text{Att}},$ (3)
In the following, we explain how we implement these two losses without using any images from past tasks.
3.3 Pseudo-Labeling for Question-only Replay
To address knowledge retention in continual VQA, we propose a pseudo-labeling approach that leverages image-question pairs formed by combining current task images with questions from past tasks stored in memory. Specifically, for each image from the current task $x^{t}$, we pair it with a question $q^{m}$ sampled from the memory. By asking past questions on new images, we prompt the model to recall and retain prior knowledge. Inspired by prior distillation-based techniques [40, 11], we employ the model from the previous task $\phi^{t-1}$ to generate answers for each new image-question pair $(x^{t},q^{m})$. These answers serve as soft pseudo-labels for the current model $\phi^{t}$, enforcing consistency with past knowledge. The pseudo-labeling loss is defined as:
$\mathcal{L}_{\text{PL}}=\mathbb{E}_{x^{t}\sim\mathcal{T}^{t}}\mathbb{E}_{q^{m}\sim\mathcal{M}}\mathcal{L}_{\text{CE}}\left[\phi^{t}(x^{t},q^{m}),\phi^{t-1}(x^{t},q^{m})\right].$ (4)
We use the network’s output as soft pseudo-labels, avoiding the application of the argmax operator. Applying argmax would compel the network to align its predictions solely with the most likely class, while soft pseudo-labeling enables a more nuanced alignment with the output distribution of the model $\phi^{t-1}$ [26, 84, 66].
Our new image-question pairs $(x^{t},q^{m})$ exposes the model to a broad set of visual-question combinations, enhancing its stability across tasks. Importantly, it allows the model to retain a broad ability to answer different types of questions, even if it cannot produce precise, correct answers. Crucially, without this pseudo-labelling regularisation, the model becomes susceptible to the out-of-answer-set problem, where overfitting to the current task’s answer space leads to incorrect responses for questions from previous tasks. For instance, after training on a color recognition task, the model might incorrectly respond with a color name when asked a counting question from a prior task.
Question Selection for QUAD. Randomly pairing images from the current task $x^{t}$ with past questions $q^{m}$ from memory can help mitigate forgetting, but it is inherently suboptimal as it often results in mismatched pairs. For instance, if the current macro-task $\mathcal{T}^{t}$ focuses on counting, random pairing could lead to questions like “How many cows are in this image?” being associated with an image of cars, which undermines the relevance and effectiveness of the training process. To address this, we introduce a question selection strategy that prioritises questions directly related to the object categories $c_{i}$ being learnt in the current visually-driven subtask $\mathcal{S}^{l}$. This targeted pairing ensures that the training process leverages meaningful and contextually relevant image-question combinations. By aligning questions with the visual context, QUAD enhances adaptation to new linguistically-driven tasks and visually-driven subtasks, leading to effective knowledge transfer.
Example: Suppose the current visually-driven subtask $\mathcal{S}^{t}$ involves learning to count cars. Thus, the question selection strategy prioritises questions from memory relevant to cars, such as “What’s the color of the car?”, ensuring alignment between the visual content and the linguistic query.
3.4 Attention Consistency Distillation
Continual VQA demands a careful balance between stability and plasticity, particularly as models must adapt to new tasks without forgetting. While pseudo-labeling offers a straightforward way to preserve prior knowledge, it provides limited stability because it only influences the network’s outputs, leaving its internal representations vulnerable to drift. Feature distillation across multiple layers [29, 77, 11, 54] provides stronger regularisation by guiding internal representations but often compromises plasticity. These methods impose rigid constraints across all layers, which is particularly problematic in our setting, where new image-question pairs $(x^{t},q^{m})$ introduced during learning the current task were not seen in previous tasks. Although these pairs are meaningful and aligned, the model has no prior experience processing them, making strict layer-wise regularisation overly restrictive. Our VQA model [85, 52] processes image features and language tokens as a unified sequence in a transformer, where self-attention naturally captures both intra-modal (language-to-language and image-to-image) and inter-modal (language-to-image) interactions. This implicit cross-attention capability enables the model to compute similarity scores between language and image tokens without requiring cross-attention layers. However, as the model is sequentially fine-tuned on new tasks, it runs the risk of suffering from self-attention drift—a gradual shift in focus away from relevant visual regions associated with previous tasks [65, 21]. This drift occurs because pseudo-labeling alone cannot prevent changes in the model’s internal representations.
To address this, we introduce an intermediate approach: attention consistency distillation. Our approach aligns the self-attention weights of the current model $\phi^{t}$ with those of the previous task model $\phi^{t-1}$, preserving both intra-modal and inter-modal relationships. This alignment acts as a lightweight but effective regularisation that maintains stability within the self-attention patterns. By distilling only the attention maps, we constrain the model to maintain its focus on relevant visual regions for answering questions without imposing strong regularisation on all layers.
Let us sample an image-question pair $(x^{t},q^{m})$ with ${}{x^{t}\sim\mathcal{T}^{t}},~{}{q^{m}\sim\mathcal{M}}$: we denote $A_{k}^{t}$ and $A_{k}^{t-1}$ as the corresponding self-attention maps in an attention head $k$ for the current and previous models, respectively, computed over the unified token sequence of both language and image features [85, 52]. These maps are generated by applying softmax to the scaled query-key product, yielding normalised distributions that represent the focus between intra-modal and inter-modal interactions. To align these distributions, we define the self-attention distillation loss $\mathcal{L}_{\text{Att}}$ using cross-entropy, encouraging the current model’s attention maps $A_{k}^{t}$ to resemble those from the previous task $A_{k}^{t-1}$:
$\begin{split}\mathcal{L}_{\text{Att}}=\mathbb{E}_{x^{t}\sim\mathcal{T}^{t}}\mathbb{E}_{q^{m}\sim\mathcal{M}}&\mathbb{E}_{k\sim\mathcal{K}_{\phi}}\\ &\mathcal{L}_{\text{CE}}\left[A_{k}^{t}(x^{t},q^{m}),A_{k}^{t-1}(x^{t},q^{m})\right],\end{split}$ (5)
where $\mathcal{K}_{\phi}$ denotes the set of all the attention heads across all the layers of $\phi$. Our approach employs cross-entropy loss over normalized attention maps, in contrast to prior works [11, 54] that rely on L1 or ReLU+L1 losses. Cross-entropy enforces alignment between attention distributions, preserving both individual token importance and structured inter-modal relationships essential for VQACL. In contrast, L1-based methods treat attention scores as independent values, imposing rigid constraints that limit adaptability to new tasks. A key advantage of cross-entropy is its consistency across all loss terms in QUAD ($\mathcal{L}_{\text{CE}},\mathcal{L}_{\text{PL}},\mathcal{L}_{\text{Att}}$), ensuring coherent optimization. Empirically, this design mitigates catastrophic forgetting.
[…]
Methods VQAv2 NExT-QA
#Mem Type Standard Test Novel Comp. Test #Mem Type Standard Test Novel Comp. Test
#Mem AP ($\uparrow$) Forget ($\downarrow$) AP ($\uparrow$) Forget ($\downarrow$) #Mem AP ($\uparrow$) Forget ($\downarrow$) AP ($\uparrow$) Forget ($\downarrow$)
Joint - None 51.64 - 51.10 - - None 35.92 - 36.24 -
Vanilla None None 14.92 30.80 11.79 27.16 None None 12.68 25.94 12.59 28.04
EWC [32] None None 15.77 30.62 12.83 28.16 None None 13.01 24.06 11.91 27.44
MAS [1] None None 20.56 11.16 23.90 6.24 None None 18.04 10.07 21.12 10.09
[…]
4 Experiments
4.1 Experimental setup
Implementation Details. To ensure a fair comparison, we adopt the protocol of [85] for both feature extraction and training across datasets. For the visual embeddings, we use a Faster R-CNN [58] trained on the Visual Genome dataset [34] to obtain 36 region-based features per image in the VQAv2 dataset. For videos in the NExT-QA dataset, we extract clip-level motion features using an inflated 3D ResNeXt-101 [25], setting $n=16$ regions per clip. These visual features are adapted through a two-layer MLP with GELU activation, preparing them as inputs to the transformer model. Our transformer backbone, based on T5 [55], consists of 12 blocks for both encoder and decoder modules, each containing 12 attention heads. The embedding dimension $d$ is tailored to the task-specific requirements. Training is conducted for 3 epochs per task, with a batch size of 80. We utilize the Adam optimizer [31] with an initial learning rate of $10^{-4}$. $\lambda$ is set to $0.5$ in all experiments. All implementations are based on PyTorch [53]. Unlike VQACL [85], QUAD does not create or store any prototypes for questions or visual objects.
Evaluation Metrics. We utilise two established metrics for continual learning [85, 4, 45]: final average performance ($AP$), and average forgetting ($Forget$). The $AP$ metric reflects the model’s overall performance across all learnt tasks, highlighting its ability to consistently acquire new tasks. Let $a_{i,j}$ represent the performance of the model on task $\mathcal{T}^{i}$ after it has completed learning this task $\mathcal{T}^{j}$. Then, $AP$ is calculated as: $AP\!=\!\frac{1}{T}\sum_{t=1}^{T}a_{t,T}.$ Additionally, the $Forget$ metric quantifies the performance loss in previous tasks as new tasks are learnt and is computed as: $Forget\!=\!\frac{1}{T-1}\sum_{t=1}^{T-1}\max_{z\in\{t,\ldots,T-1\}}(a_{t,z}-a_{t,T}).$ To ensure a fair comparison, we calculate $a_{i,j}$ in the NExT-QA dataset as per the approach in [85, 75], utilising Wu-Palmer Similarity (WUPS) [74] to assess the quality of the generated answers. For the VQAv2 dataset, following the method outlined in [85], we use the percentage of correctly answered questions as the value for $a_{i,j}$.
Baselines. To rigorously evaluate the performance of our method, we conduct a comprehensive analysis encompassing five established and state-of-the-art continual learning methods. These methods include two regularisation-based approaches: Elastic Weight Consolidation (EWC) [33] and Memory Aware Synapses (MAS) [1], as well as three rehearsal-based approaches: Experience Replay (ER) [7], Dark Experience Replay (DER) [6], Virtual Sample (VS) [67], and VQACL [85] (see details in Appendix). To contextualise the performance of these methods, we establish two performance boundaries: a lower bound (Vanilla) that performs simple gradient updates without any forgetting mitigation strategies, and an upper bound (Joint) that trains on all tasks simultaneously, providing an upper bound performance metric. Following the evaluation protocol established in VQACL [85], we employ two key evaluation strategies: standard testing, which evaluates the models’ performance on previously encountered task types, assessing their ability to retain learnt knowledge; and novel composition testing, which challenges the models with previously unseen combinations of visual and linguistic elements, probing their capacity for compositional generalization. This dual-pronged evaluation strategy, adopted from the original VQACL work, provides a nuanced understanding of each method’s efficacy in balancing knowledge retention and adaptive learning in the VQACL.
Table 2: Fine-grained VQA performance AP (%) on the Novel and Seen skill-concept compositions of VQAv2 and NExT-QA. +$\Delta$ denotes the improvement of our method over the state of the art. Best results are in bold, second-best are underlined.
Dataset Method Memory Group-1 Group-2 Group-3 Group-4 Group-5 Avg
Novel Seen Novel Seen Novel Seen Novel Seen Novel Seen Novel Seen
[…]
4.2 Main results
Performance Analysis On Standard Setting. Tab. 1 shows a detailed comparison of continual learning approaches in the VQACL setting, with our proposed QUAD achieving superior performance across both standard and novel composition tests. QUAD consistently outperforms other methods in AP and forgetting demonstrating robust knowledge retention. On VQAv2, QUAD attains an AP of 39.25% in standard testing, outperforming the best rehearsal-based approach (VQACL) by 1.79%. Similarly, in NExT-QA, QUAD achieves an AP of 31.70%, exceeding competing methods by 0.84% to 4.73%. The forgetting rates for QUAD are also the lowest, with 4.91% and 2.91% for VQAv2 and NExT-QA, respectively, underscoring its capacity for stable knowledge retention.
In novel composition testing, QUAD demonstrates strong generalization, achieving top AP scores of 40.00% on VQAv2 and 33.85% on NExT-QA. Notably, despite not storing images, our approach maintains high performance in novel settings, with a minimal performance gap between standard and novel composition tests (0.75% for VQAv2 and 0.64% for NExT-QA). This narrow gap indicates that distilling knowledge using only seen questions effectively reinforces visual-linguistic associations, enabling the model to generalise well even in unfamiliar contexts. This performance underscores QUAD’s capacity to construct robust, adaptable representations, showing that question-only distillation can successfully support compositional reasoning without the need for visual memory.
Performance Analysis of Novel Composition Testing. Tab. 2 provides a detailed comparison of model performance on novel and seen skill-concept compositions across VQAv2 and NExT-QA datasets. QUAD consistently surpasses previous approaches. On VQAv2, QUAD shows significant gains over VQACL, with an average improvement of 4.60% on novel compositions and 3.81% on seen compositions. This result indicates enhanced compositional generalisation, as evidenced by the smaller gap between novel and seen performance compared to other methods. For NExT-QA, QUAD maintains competitive performance, with slight gains over VQACL on seen groups (average +0.27%) in a dataset that presents unique challenges, such as temporal and causal reasoning. For such tasks, storing images may be necessary to maintain a comprehensive task understanding. These consistent gains highlight QUAD’s robustness in compositional reasoning, validating its effectiveness for continual VQA.
5 Ablation Study and Analysis
Figure 3: Plasticity/Stability analysis on VQAv2. Each matrix shows the performance of a model trained on tasks (rows) and evaluated on tasks sequentially (columns). The diagonal (highlighted in orange) represents in-domain performance, while off-diagonal elements indicate cross-domain generalization. Higher values (darker colors) suggest better retention and plasticity. The progression from ‘Sequential finetuning’, to $\mathcal{L}_{PL}$ and then to our full method, QUAD, highlights the improvement in retaining knowledge across sequential tasks.
1
Distillation Components. Tab. 3 shows the impact of PL distillation and attention distillation in QUAD. We evaluate three setups: (1) PL distillation alone, (2) attention distillation alone, and (3) the combination of both.
Using PL distillation alone achieves moderate performance, with AP of 30.72% on VQAv2 and 29.04% on NExT-QA, indicating effective feature alignment across tasks. In contrast, attention distillation alone yields lower AP scores (13.34% on VQAv2 and 13.24% on NExT-QA) with high forgetting scores (32.08% on VQAv2 and 24.56% on NExT-QA), suggesting limited task adaptation. Combining PL and attention distillation achieves the best results, with AP scores of 39.25% on VQAv2 and 31.70% on NExT-QA, and the lowest forgetting rates.
2
Attention Distillation. Tab. 4 demonstrates the effectiveness of our proposed $\mathcal{L}_{\text{Att}}$ within QUAD. Unlike prior methods such as Attn-dist (L1) and Asym-Attn [54], which impose L1 or ReLU+L1 losses on raw attention scores, our approach applies cross-entropy over normalized attention maps. QUAD outperforms previous methods, achieving an AP of 39.25% and Forgetting of 4.91% on VQAv2, and an AP of 31.70% with Forgetting of 2.91% on NExT-QA. These results highlight the importance of preserving distributional consistency in attention maps and avoiding unstructured alignment.
Table 3: Ablation study of QUAD components on VQAv2 and NExT-QA.
$\mathcal{L}_{\text{PL}}$ $\mathcal{L}_{\text{Att}}$ Memory VQAv2 NExT-QA
Type AP (↑) Forget (↓) AP (↑) Forget (↓)
[…]
Table 4: Comparison of attention distillation methods.
Method VQAv2 NExT-QA
AP (↑) Forget (↓) AP (↑) Forget (↓)
$\mathcal{L}_{\text{PL}}$ + Attn-dist (L1) 34.56 7.91 30.14 5.78
$\mathcal{L}_{\text{PL}}$ + Asym-Attn [54] 38.15 5.57 31.18 4.13
QUAD (Ours) 39.25 4.91 31.70 2.91
3
Analysing Plasticity/Stability. Fig. 3 illustrates the impact of different strategies on plasticity and forgetting in the VQAv2 dataset. The sequential fine-tuning matrix (left) reveals the impact of learning without rehearsal, showing uniformly low off-diagonal values—indicating the model’s failure to answer questions from previous tasks. This phenomenon, which we term the out-of-answer-set problem, occurs when the model overfits to the current task’s answer space, preventing it from correctly responding to questions from earlier tasks. The $\mathcal{L}_{\text{PL}}$ matrix (center) mitigates forgetting through pseudo-label distillation, significantly improving cross-domain retention in tasks such as “commonsense” and “count.” However, it remains less effective in complex reasoning tasks like “causal” and “subcategory.” Our full method, QUAD (right), achieves the best overall performance by integrating pseudo-label and attention consistency distillation, striking a balance between stability and plasticity. QUAD consistently maintains high values both on and off the diagonal, demonstrating superior retention of prior knowledge and adaptability to new tasks. Attention consistency can be assessed by comparing the right and middle subplots. For instance, after training with QUAD, the model retains more knowledge on the ‘judge’ task (62.6%) compared to pseudo-labeling alone (35.0%).
4
Sensitivity to Memory Size. Fig. 4 shows a sensitivity analysis of CL methods with respect to memory size. QUAD consistently outperforms baseline methods (ER, DER, VS, VQACL) across all memory sizes, demonstrating the effectiveness of our distillation strategy. On VQA-v2, QUAD exhibits strong scalability, with AP increasing steadily from 1000 to 5000 samples. In constrained storage scenarios, QUAD maintains competitive performance by leveraging question-only distillation. On NExT-QA, QUAD outperforms other methods, though with smaller gains than on VQAv2 due to the increased task complexity and reliance on visual information.
Figure 4: Sensitivity analysis to memory size. Our method, QUAD, consistently achieves higher AP than baselines, demonstrating strong scalability on VQAv2 and stable performance on NExT-QA, especially as memory size increases.
5
Effectiveness of Object-Matched Question Selection. Fig. 5 demonstrates the benefits of object-matched question selection over random selection. We analyze its impact across memory sizes on VQAv2 and NExT-QA, where it consistently improves AP as memory grows. Aligning selected questions with object categories enhances knowledge transfer, leading to better adaptation to new tasks.
Figure 5: Question selection strategy in QUAD. Figure shows AP for random (blue) and object-matched (purple) question selection across memory sizes on VQAv2, and NExT-QA. Object-matched selection consistently outperforms random selection.
6 Conclusion
In this work, we introduced QUestion-only replay with Attention Distillation (QUAD), a novel distillation-based approach for continual VQA. Unlike conventional methods that store both images and questions, QUAD addresses storage and privacy concerns by retaining only past task questions. This design enables effective regularization without storing sensitive visual data, making it highly practical for privacy-conscious applications.
Comprehensive evaluations on VQAv2 and NExT-QA show that QUAD consistently outperforms both memory-free and memory-rehearsal methods, achieving state-of-the-art generalization and knowledge retention. These results validate question-only replay as an efficient and scalable solution for continual VQA. Future work could explore dynamically generated questions to eliminate memory storage entirely, further enhancing adaptability in real-world multimodal scenarios.
Acknowledgements. This paper has been supported by the French National Research Agency (ANR) in the framework of its JCJC, and was funded by the European Union’s Horizon Europe research and innovation program under grant agreement No. 101120237 (ELIAS). Furthermore, this research was partially funded by Hi!PARIS Center on Data Analytics and Artificial Intelligence. This work was granted access to the HPC resources of IDRIS under the allocation AD011013860 made by GENCI.
Supplementary Material
In this supplementary material, we provide additional details on the experimental results presented in the main paper, along with further empirical evaluations and discussions.
Section 7 presents an ethics statement regarding our work. Section 8 discusses the limitations of our proposed method, QUAD. Section 9 reports the computational and memory footprint analysis of QUAD. Section 10 and Section 11 describe the datasets, task sequences, and evaluation structure. Section 12 provides analysis on plasticity-stability for NExT-QA dataset. Lastly, Section 13 explains the baselines used for comparison.
7 Ethics Statement
Our method, QUAD, is designed to improve continual learning in Visual Question Answering (VQACL) while maintaining generalization and privacy through distilaltion using questions-only. We do not foresee any negative societal impact from this work, as it does not involve the generation of harmful or biased data. However, like any machine learning system, there remains a potential risk if it is applied unethically or without proper oversight. QUAD’s design includes mechanisms to enhance privacy, reducing the storage of sensitive visual data. Despite this, its applicability beyond the specific datasets and tasks used in our experiments remains to be thoroughly tested, and we caution against the unconsidered deployment of the method in sensitive applications without further validation.
8 Limitations of QUAD
While QUAD effectively reduces storage requirements and enhances privacy by eliminating the need to store images, it may be suboptimal for tasks that heavily rely on detailed visual or spatial reasoning. Certain VQA tasks, such as object classification, fine-grained attribute recognition, or spatial relationships, inherently require access to visual information to retain critical knowledge from previous tasks. For instance, as shown in Fig. 3, QUAD struggles to maintain performance on the ‘type’ task in VQAv2, which depends on visual cues, whereas it performs well on conceptually driven tasks like ‘commonsense’ reasoning.
Our findings suggest that question-only replay is particularly well-suited for constrained scenarios where privacy and storage efficiency are primary concerns. However, in settings where high fidelity in visual reasoning is essential, storing a subset of representative images may be necessary to preserve task-specific knowledge and improve overall performance. Future work could explore hybrid approaches that selectively retain visual information while leveraging question-based replay, striking a balance between efficiency and task-specific retention.
Furthermore, QUAD prevents storing original sensitive visual information, aligning with GDPR constraints, which permit data storage only when strictly necessary for the task. However, our approach specifically addresses storage-related privacy concerns and does not guarantee protection against attacks such as inversion attacks [12, 82].
9 Computational Analysis
Efficient memory and storage management is crucial for continual VQA, where scalability is a key challenge. This section analyzes storage requirements, computational complexity, and GPU memory usage of our text-only replay approach compared to image-based methods. By storing only past task questions, we significantly reduce storage complexity from $\mathcal{O}(N\cdot(I+L_{q}+L_{a}))$ to $\mathcal{O}(N\cdot L_{q})$, where $N$ is the number of stored samples, $I$ is the image size, and $L_{q}$ and $L_{a}$ represent the question and answer lengths in bits.
In terms of GPU memory usage, question-only replay has a minimal impact since the number of processed input pairs remains the same. The primary reduction stems from loading fewer images, but this accounts for less than 5% of the total memory footprint, which is dominated by gradients, weights, and activations. This makes our approach particularly appealing in scenarios where storage is constrained but GPU memory availability remains a concern.
From a computational complexity perspective, our method does not introduce any additional overhead. The computational cost remains unchanged when processing images from past or current tasks. The forward and backward passes are identical, ensuring that our approach maintains the same efficiency while significantly improving storage scalability.
This analysis validates our design choices, demonstrating that question-only replay can achieve competitive performance while substantially reducing storage requirements. This efficiency makes it highly scalable and practical for real-world deployment.
10 Detailed Description of the VQACL Setting
This section provides a detailed overview of the Visual Question Answering Continual Learning (VQACL) setting, as introduced by [85]. The VQACL setting is designed to test a model’s ability to generalise and retain knowledge across a sequence of tasks involving both visual and linguistic modalities, with a particular focus on compositional generalisation and knowledge retention.
The VQACL setting is organised into a two-level hierarchy of tasks that challenge both the visual and linguistic capabilities of the model.
Table 5: Linguistic-driven task statistics of VQA v2 in the VQACL setting. Stan. Test denotes the standard test set.
Task Train Val Stan. Test Examples
Recognition 131,478 5,579 5,628 What is on the floor? What does the sign say?
Location 12,580 611 611 Where is the giraffe? Where are the people standing?
Judge 160,179 7,126 7,194 Is the baby playing ball? Are the windows big?
Commonsense 25,211 1,114 1,100 Do the elephants have tusks? Do the dogs know how to swim?
Count 62,156 2,651 2,658 How many beds? How many seats are there?
Action 33,633 1,498 1,373 Are they drinking wine? Is the person flying?
Color 50,872 2,322 2,192 What color is the bedspread? What color are the gym shoes?
Type 23,932 1,119 1,089 What type of building is this? What type of animal is shown?
Subcategory 31,594 1,477 1,416 What brand is the umbrella? What brand are his shoes?
Causal 5,868 231 200 Why does he have glasses on? Why is the dog jumping?
Table 6: Linguistic-driven task statistics of NExT-QA in the VQACL setting. Stan. Test denotes the standard test set. CW: CausalWhy; TN: TemporalNext; TC: TemporalCurrent; DL: DescriptiveLocation; DB: DescriptiveBinary; DC: DescriptiveCount; DO: DescriptiveOther; CH: CausalHow.
Task Train Val Stan. Test Examples
CW 13,552 1,928 3,333 Why is the lady sitting down? Why is the baby’s hair wet?
TN 5,685 895 1,399 What does the baby do after picking up the toy? What did the lady do after adjusting the shirt?
TC 4,797 663 1,165 What event is happening? What sport is the man doing?
DL 1,942 295 482 Where are the two people dancing? Where is this video taken?
DB 2,928 277 495 Is the baby able to walk? Does the girl cry?
DC 1,378 192 365 How many babies are there? How many dogs are there?
DO 2,549 356 672 What season is this? What does the man use to stir the food in the pan?
CH 4,400 683 1,174 How did the singer project her voice? How did the boy in the box move forward?
Table 7: Detailed information about the five object groups in VQA v2.
Task Objects
Group 1 hot dog, fork, orange, snowboard, potted plant, person, toilet, laptop, surfboard, bench, bus, dog, knife, pizza, handbag, bicycle
Group 2 horse, cell phone, elephant, boat, zebra, apple, stop sign, microwave, spoon, cup, skateboard, tie, umbrella, sandwich, bear
Group 3 donut, truck, frisbee, giraffe, dining table, motorcycle, parking meter, car, oven, airplane, bed, sheep, baseball bat
Group 4 skis, baseball glove, tennis racket, tv, traffic light, kite, cake, keyboard, bottle, remote, bird, carrot
Group 5 suitcase, couch, broccoli, cow, fire hydrant, chair, mouse, cat, banana, wine glass, backpack, bowl, sports ball, train
Table 8: Detailed information about the five object groups in NExT-QA.
Task Objects
Group 1 bicycle, camel, bat, microwave, snake, sofa, traffic light, hamster/rat, chicken, oven, stop sign, vegetables, skateboard, bird, toilet, racket
Group 2 crab, camera, lion, ball/sports ball, crocodile, screen/monitor, baby walker, cat, squirrel, frisbee, cattle/cow, sheep/goat, adult, scooter, electric fan, stool
Group 3 piano, watercraft, kangaroo, train, fruits, pig, suitcase, bear, tiger, bench, elephant, motorcycle, horse, snowboard, surfboard, handbag
Group 4 ski, stingray, antelope, toy, child, duck, guitar, dish, fish, cake, turtle, leopard, laptop, panda, table, cup
Group 5 penguin, faucet, car, bottle, bus/truck, aircraft, baby, bread, baby seat, cellphone, sink, rabbit, backpack, chair, dog, refrigerator
  • Linguistically-Driven Tasks. At the higher level, the VQACL setting comprises a series of linguistically-driven tasks, denoted as ${\mathcal{T}^{1},\dots,\mathcal{T}^{T}}$, where $T$ represents the total number of tasks. Each task focuses on a specific reasoning skill, such as counting or color identification, and is characterized by a particular type of question. For example, a task focused on counting might involve questions beginning with ”How many” or ”What number”. In our experiments, the VQAv2 dataset consists of $T\!=\!10$ such tasks, while the NExT-QA dataset includes $T\!=\!8$ tasks.
  • Visually-Driven Subtasks. Nested within each linguistically-driven task are a series of visually-driven subtasks ${\mathcal{S}_{1}^{t},\dots,\mathcal{S}_{K}^{t}}$. Each visually-driven subtask is associated with a specific object group $G_{k}$, formed by partitioning the total set of object classes ${\{c_{i}}\}_{i=1}^{C}$ into $K$ groups. These groups are then randomly assigned to different subtasks within each linguistic-driven task. In our implementation, both the VQAv2 and NExT-QA datasets are divided into $K\!=\!5$ visual subtasks, covering a total of $C\!=\!80$ object classes, following the categorization used in the COCO dataset [41].
  • Novel Composition Testing. The VQACL setting also includes a novel composition testing process, designed to evaluate the model’s compositional generalization abilities—its capacity to apply learned concepts to new combinations of objects and questions.
Training and Testing Procedure. During training, the model is exposed to a subset of the visual-driven subtasks within each linguistically-driven task. Specifically, one visual-driven subtask $\mathcal{S}_{k}^{v}$ is randomly excluded from the training phase for each linguistic-driven task. This excluded subtask is reserved for testing and serves as a novel composition, where the model must answer questions about unseen combinations of objects and reasoning skills.
Cross-Validation and Fair Testing. To ensure a fair evaluation of the model’s generalization capabilities, the VQACL setting employs a $K$-fold object-independent cross-validation process. This involves repeating the training and testing procedure $K$ times, each time excluding a different visual-driven subtask. This ensures that the model encounters all object classes across different folds, thereby providing a comprehensive assessment of its ability to generalize to new combinations of objects and tasks.
Continual Learning Challenges. The VQACL setting presents a significant challenge for continual learning models, requiring them to balance the retention of knowledge from previously learnt tasks (stability) with the ability to adapt to new, continually arriving tasks (plasticity). By structuring tasks to involve both new and previously encountered concepts, the VQACL setting effectively tests the model’s ability to minimize catastrophic forgetting while enabling knowledge transfer across tasks.
11 Details of Evaluation Datasets
In this section, we provide a detailed overview of the two datasets used in our evaluation: VQA v2 and NExT-QA. Each dataset has been carefully structured into different tasks, which are used to evaluate the performance of our continual learning models.
We summarize the statistics of each dataset, focusing on both linguistic and object-related tasks. Tables 5 and 6 (previously described) present the linguistic-driven task breakdown, including categories such as Recognition, Commonsense, Count, and others.
Additionally, we grouped the objects in each dataset into five distinct object groups to facilitate better understanding and comparison of the models’ object recognition capabilities. Tables 7 and 8 offer a detailed breakdown of the objects associated with each group in VQA v2 and NExT-QA, respectively. This categorization will aid in analyzing how the models perform across different object categories.
These two datasets, each structured uniquely in terms of linguistic tasks and object types, allow us to rigorously assess the models in varied real-world scenarios. Together, these benchmarks enable a comprehensive evaluation of the continual learning approaches proposed in this work.
12 Extended Analysis of Plasticity/Stability Trade-Off
Fig.6 compares the impact of three continual learning strategies on performance across tasks in the NExT-QA dataset. The sequential finetuning baseline (left) demonstrates severe forgetting, with consistently low off-diagonal values. Specifically, tasks like temporal reasoning (TN and TC) exhibit the worst performance, as these tasks require advanced reasoning over time sequences, which is inherently challenging for the model.
Introducing pseudo-label distillation through $\mathcal{L}_{\text{PL}}$ (center) mitigates the issue of forgetting by enforcing output consistency with the previous model. This results in improved cross-domain retention, particularly in easier tasks like ‘DB’ and ‘DL’. However, its performance on complex tasks such as ”DO” (Descriptive Others) and ‘CH’ (Causal How) remains suboptimal, as these tasks require the model to maintain intricate visual-linguistic relationships, which $\mathcal{L}_{\text{PL}}$ alone struggles to address.
Our method, QUAD (right), achieves the highest overall performance by combining pseudo-labeling with attention consistency distillation. This dual mechanism effectively balances stability and plasticity, as evidenced by the consistently high diagonal values and substantial improvements in off-diagonal cross-domain generalization. Notably, QUAD performs significantly better on retraining prior knowledge (row 6, 7). The results underscore the strength of QUAD in preserving visual-linguistic associations and mitigating the out-of-answer-set problem across tasks in NExT-QA.
Figure 6: Comparison of feature distillation methods on NExT-QA. Each matrix shows the performance of a model trained on tasks (rows) and evaluated on tasks (columns). The diagonal (highlighted in orange) represents in-domain performance, while off-diagonal elements show cross-domain generalization. Higher values (darker colors) indicate better performance.
13 Continual Learning Methods
We assess and benchmark five prominent continual learning methods, encompassing two regularization techniques (EWC [33], MAS [1]) and three rehearsal-based methods (ER [7], DER [6], VS [67], and VQACL[85]). To ensure a consistent evaluation, all methods are implemented using their official codebases and integrated into the same transformer backbone as described in Section 5.1.
EWC [33] is a regularization method designed to preserve knowledge of prior tasks by selectively reducing updates on critical parameters. This is achieved by leveraging the Fisher Information Matrix, which quantifies the importance of parameters and incorporates an auxiliary L2 loss between significant parameters from old and new tasks.
MAS [1] similarly applies regularization, aiming to prevent significant changes to parameters vital for previous tasks by introducing an L2 loss. In contrast to EWC, MAS measures the sensitivity of the output with respect to parameter perturbations to estimate parameter importance.
ER [7] is a rehearsal method that utilizes a fixed-size memory buffer, where visited examples are stored and randomly sampled for retraining. In line with our approach, the memory size for ER is fixed at 5,000 for VQA v2 and 500 for NExT-QA. Given its simplicity and effectiveness, ER serves as the baseline for our proposed method.
DER [6] is another rehearsal technique that employs reservoir sampling to manage memory, ensuring every visited sample has an equal chance of being stored. DER also incorporates a dark knowledge distillation strategy, which aims to align the network’s outputs with logits recorded during training, thus encouraging consistency in responses to prior examples. In our experiments, DER also utilizes memory sizes of 5,000 for VQA v2 and 500 for NExT-QA.
VS [67] is a rehearsal-based method that emphasizes feature consistency between current and past data. To address forgetting, VS introduces two losses: a neighbor-session model coherence loss and an inter-session data coherence loss. For more details, we refer readers to Wan et al. [67]. The memory size for VS is similarly set to 5,000 for VQA v2 and 500 for NExT-QA.
VQACL [85] represents a rehearsal-based approach, incorporating a prototype module to learn both task-specific and invariant features, facilitating robust and generalizable representations for VQA tasks.